determine whether the set s={(a,b,c) ; ab=1} is a subspace of R3


1
Expert's answer
2021-11-29T19:19:23-0500



If "\begin{pmatrix}
 0 \\
 0\\
0
\end{pmatrix}" "\in\>V" then given "ab=1," then "(0)(0)\neq1"


"\therefore \begin{pmatrix}
 0 \\
 0\\0
\end{pmatrix}" is not V "\implies" V is empty



2). Let "x,y\in" V then


"x=\begin{pmatrix}
 x_1 \\
 x_2 \\
x_3
\end{pmatrix}" and "y=\begin{pmatrix}
 y_1 \\
 y_2 \\
y_3
\end{pmatrix}" "x_iy_i\in\R"



"x_1x_2=1" and "y_1y_2=1"

"x_2=\frac{1}{x_{1}}" "y_2=\frac{1}{y_1}"


"x+y= \begin{pmatrix}
 x_1+y_1 \\
 x_2+y_2 \\
x_3+y_3
\end{pmatrix}"



"(x_2+y_2)(x_1+y_1)=1\implies(x_2+y_2)=\frac{1}{x_1+y_1}"



"x_2+y_2=\frac{1}{x_1}+\frac{1}{y_1}\neq\frac{1}{x_1+y_1}"



S is not closed under addition


3) Take "x\in" V and "\alpha\in\R" then


"x=\begin{pmatrix}
 x_1 \\
 x_2\\ 
x_3
\end{pmatrix}" for "x_i \in\R"



"(x_1)(x_2)=1\implies\alpha\begin{Bmatrix}
 (x_1)(x_2) 
 
\end{Bmatrix}=\alpha(1)"



"\alpha\>x=\begin{pmatrix}
 \alpha\>x_1 \\
 \alpha\>x_2\\
\alpha\>x_3
\end{pmatrix}"


"\therefore(\alpha\>x_2)(\alpha\>x_1)=\alpha^2(x_1)(x_2)\ne\alpha(x_1)(x_2)"



S is not closed under scalar multiplication


S is not a subspace of "\R^3"